Expand and simplify each function. Evaluate at f(-2). a. f(x) = (x + 1)(x - 2) (x + 3) b. f(x) = (2x ^ 2 - 3)(x ^ 3 + 1)


1
Expert's answer
2021-02-24T06:32:31-0500

"\left(a\right).\:\left(x+1\right)\left(x-2\right)\left(x+3\right)"

"\left(x\left(x-2\right)+1\left(x-2\right)\left(x+3\right)\right)"

"\left(x^2-2x+x+2\right)\left(x+3\right)"

"\left(x^2-x+2\right)\left(x+3\right)"

"x^3-x^2+2x+3x^2-3x+2\:"

"x^3+2x^2-x+2"

"at\:f\left(-2\right)=-2:"

"=\left(-2\right)^3+2\left(-2\right)^2-\left(-2\right)+2\:"

"=-8+8+2+2\:"

"=4"


"\left(b\right).\:\left(2x^2-3\right)\left(x^3+1\right)"

"2x^2\left(x^3+1\right)-3\left(x^3+1\right)"

"2x^5+2x^2-3x^3-3"

"2x^5+3x^3+2x^2-3"

"at\:f\left(x\right)=-2"

"=2\left(-2\right)^5+3\left(-2\right)^3+2\left(-2\right)^2-3"

"=-64-24+8-3"

"=-83"